# library.bib

### v1

@@ Abstract
A planar stuffed map is an embedding of a graph into the 2-sphere $S^{2}$, considered up to orientation-preserving homeomorphisms, such that the complement of the graph is a collection of disjoint topologically connected components that are each homeomorphic to $S^{2}$ with multiple boundaries. This is a generalization of planar maps whose complement of the graph is a collection of disjoint topologically connected components that are each homeomorphic to a disc. In this work, we construct a bijection between certain bipartite planar stuffed maps and collections of integer-labelled trees connected by hyperedges such that they form a hypertree, called hypermobiles. This bijection directly generalizes the Bouttier-Di Franceso-Guitter bijection between bipartite planar maps and mobiles. We show that the generating functions of these trees of mobiles satisfy both an algebraic equation, generalizing the case of ordinary planar maps, and a new functional equation. As an example, we explicitly enumerate a class of stuffed quadrangulations.
@@ 1 Introduction
Planar maps were first studied and enumerated by Tutte in the 1960’s [tutte1968enumeration] and have since appeared in statistical mechanics, quantum gravity, and probability theory. For a detailed history see [schaeffer2015planar]. One key insight discovered was that planar maps are in bijection with decorated trees [cori1981planar, schaeffer1998conjugaison] from which enumerative formulae could be derived. Of particular relevance to this work is the bijection of Bouttier-Di Franceso-Guitter (BDFG) [bouttier2004planar], which is between planar maps and plane trees with integer-labelled vertices called mobiles. Such bijections have since been generalized beyond planar maps [chapuy2009bijection, miermont2009tessellations]. In addition to enumeration, these bijections gave insight into the discrete geometry of maps which has found applications in theoretical physics and probability theory [bouttier2003geodesic, le2010scaling].
@@
As the reader will see, stuffed maps have many interesting combinatorial properties, and their enumeration is much more complicated than ordinary maps. Stuffed maps were first introduced in [borot2014formal], where it was shown that stuffed maps of any genus with boundaries could be recursively enumerated using a process called blobbed topological recursion. This work was largely motivated by the study of multi-tracial matrix models. As models of quantum gravity, the multi-tracial terms in the potential of such matrix models contribute “wormholes" in the two-dimensional universes of their Feynman diagrams i.e. stuffed maps. In particular, an interior of a 2-cell with multiple boundaries corresponds to a wormhole and a path-connected component corresponds to universes in the physics literature [das1990new, ambjorn2001lorentzian]. Such matrix models are also of interest in what is known as the duality of Liouville quantum gravity [klebanov1995non, duplantier2009duality, rhodes2014gaussian]. Other applications include the $O(n)$ model which involves ordinary maps decorated with self-avoiding loop configurations which can be shown to be equivalent to stuffed maps as well as maps with intersecting loops.
We hope that this paper will lay the foundations for studying the geometric and statistical properties of stuffed maps. Stuffed maps have yet to be fully studied from the perspective of random maps. Recently, much work has been done to connect the Gromov-Hausdorf limit of random metric spaces to random surfaces from two-dimensional quantum gravity. Researchers are looking for new random surfaces, and many types of maps have been shown to converge to such spaces [le2019brownian]. It is believed, based on critical exponent analysis, that stuffed maps will not fall into the same class as ordinary maps, and hence might be related to new types of random surfaces. One can enumerate stuffed maps by computing formal multi-tracial matrix integrals, and in the physics literature such integrals are conjectured to have a continuum limit with "baby-universes" [korchemsky1992matrix, ambjorn1993baby]. Certain coloured stuffed maps appear also in the study of tensor models [bonzom2017colored]. Establishing bijections with trees could be used as a first step in establishing a rigorous construction of a random surfaces with such features, as has been done for many types of ordinary maps [le2010scaling, le2013uniqueness, bettinelli2017compact, gwynne2017scaling].
@@ 2.2 Mobiles
  • For any white vertex labelled $n$ that is attached to a black vertex, any adjacent clockwise white vertex attached to the same black vertex must be labelled at least $n-1$.
@@
The BDFG bijeciton is between pointed rooted bipartite planar maps and rooted mobiles [bouttier2004planar], which generalized the bijection in [schaeffer1998conjugaison] between quadrangulations and labelled trees.
@@
    For every corner of a white vertex in $m$, if the label of the vertex is $n_{\text{min}}(m)$ then connect it to the white vertex added for that mobile in the previous step. Otherwise, if the label is $n\not=n_{\text{min}}(m)$, then connect it to the next white vertex labelled $n-1$ in a clockwise direction around the contour. If this corner is the corner to the left of a root, the new edge is rooted pointing away for the corner.
@@ 2.3 Hypermobiles
  • A set of mobiles $\mathcal{M}$ with a unique rooted mobile with the lowest labelled vertex, called the origin mobile.
@@ 2.4 Branch-pointed stuffed maps
As mentioned in the Introduction, one motivation to study stuffed maps is to establish convergence, in some sense, as a random discrete metric space to a random continuous metric space. All methods known to the author require a bijection to some path-connected tree-like object. However, one distinguishing feature of stuffed maps is that they may have distinct path-connected components connected by branches, which makes it not always possible to label vertices with their graph distances from the source vertex. To facilitate such bijections, we will slightly change the type of stuffed map studied in this work by decorating them and restricting their 2-cells. Define a branch-pointed stuffed (BPS) map as a pointed stuffed map whose faces come from polygons of any size and multi-boundary 2-cells where each boundary has a distinguished vertex called a spurious point. The source cannot coincide with a spurious point. One can define the following set of BMS maps analogously to our definition of $\mathbb{M}_{\ell}(S;v)$ for stuffed maps: denote it $\overline{\mathbb{M}}_{\ell}(S;v)$.
Working with rooted BMS maps allows one to give each vertex a positive integer label, which is key to the bijection. Note that on the level of enumeration, computing the generating function of BMS maps suffices to enumerate stuffed maps and vice versa. For more details, see the Remark in Section 3.1.
@@ 2.5 A bijection with hypermobiles
In this section we will construct and prove the main bijection between planar BMS maps and hypermobiles. First, consider the mapping $\Phi$ that takes a rooted pointed BMS map $M\in\overline{\mathbb{M}}^{\circ}(S;v)$ to a hypergraph $\Phi(M)$, constructed as follows:
@@
Figure 4: Stages of the construction of a hypermobile from the BMS map seen in Figure 1 (b). Sub-figure (a) is a BMS map whose vertices are labelled with the distance of the source and whose edges next to gate vertices are connected by hyperedges. In sub-figure (b), edges are drawn according to the same rules as the BDFG bijection on each component. Sub-figure (c) is the resulting hypermobile after removing all the original edges and branches, and shifting labels.
@@
    For every corner of a white vertex in $m$, then connect it to the next vertex in a clockwise direction with $n-1$ in the contour. If this corner is the corner to the left of a root, the new edge is rooted pointing away for the corner.
@@
    Remove all the original edges and black vertices of $m$. Replace each hyperedge with a 2-cell whose boundaries correspond to the hyperedges’ vertices. The minimum labelled vertex on the origin mobile is taken to be the origin. Spurious points are taken to the vertex in a counterclockwise direction along each boundary face of a gate vertex.
@@
The mapping $\Phi$ is equal to $\Psi^{-1}$ and is a bijection between rooted and pointed BMS maps $\overline{\mathbb{M}}^{\circ}(S;v)$ and hypermobiles $\mathbb{T}(S;v)$.
@@
First, we will show that $\Phi$ maps into $\mathbb{T}(S;v)$. Let $M\in\overline{\mathbb{M}}^{\circ}(S;v)$. It is clear that $\Phi(M)$ is a labelled hypergraph. On each component one has applied the BDFG bijection, so they correspond to mobiles. The resulting graph is a collection of mobiles connected by the occasional hyperedges. Consider a host graph $G$ formed from $\Phi(M)$ by replacing each hyperedge with a linear sequence of $n-1$ red edges connecting all black vertices in that hyperedge. If we can show that $G$ is a tree, then this suffices to show that $\Phi(M)$ is a hypertree. Suppose for purposes of contradiction that $G$ contains a loop. No loop can be formed with just ordinary edges since this would contradict that one of the mobiles is a tree. Thus, such a loop must contain at least one red edge. However, any such sequence of edges that forms a cycle would have to come from a sequence of branches and components that topologically forms a handle in $M$, which contradicts the planarity of $M$. Thus, $G$ has no cycles.
What remains is to show that $G$ is connected. By construction, recall that there are $F(M)$ many black vertices, and $V(M)-K(M)$ many white vertices. The number of edges between white and black vertices is $E(M)$, since for each existing edge precisely one edge is created. The number of red edges is $\sum_{\ell}|\ell|$, where the sum is over all distinct linear sequences $\ell$ of red edges in $G$. Note that any such sequence corresponds to the number of boundaries of the corresponding branch in $M$ minus one, i.e.
@@
where $B_{i}(M)$ denotes the number of branches in $M$ with $i$ boundaries. Recall from Section 2.1 the associated hypertree to $M$ and note that it must have a host tree. As a tree, this implies that by its construction $B(M)=K(M)-1$ and the degree sum formula can be written as $\sum_{i=2}iB_{i}(M)=2B(M)$. Putting this all together:
@@ 3.1 Generating functions
From generalizing the well-known Euler characteristic argument (see Theorem 1.2.1 of [eynard2016counting] for example) we will show that $\mathbb{M}_{\ell}(S;v)$ is a finite set, implying that $\mathcal{T}_{\ell}(t,\mathbf{t})$ is a well-defined formal series.
@@
Let $\overline{\mathcal{T}}_{\ell}(t,\mathbf{t})$ denote the generating function above but for BMS maps. The difference between the generating functions of stuffed maps and branch-rooted stuffed maps is a symmetry factor of $\ell_{1}\cdots\ell_{k}$ for every 2-cell with boundaries of lengths $\ell_{1},...,\ell_{k}$ and a factor of $t$ for each boundary, i.e.
$\mathcal{T}_{\ell}\left(t,...,\frac{t_{\ell_{1},...\ell_{k}}}{t^{k}\ell_{1}\cdots\ell_{k}},...\right)=\sum_{v=1}^{\infty}t^{v}\sum_{\gamma\in\mathbb{M}_{\ell}(S;v)}\prod_{i=1}t_{i}^{n_{i}(M)}\prod_{k=2}\prod_{\ell_{1},...\ell_{k}=1}\frac{t_{\ell_{1},...\ell_{k}}^{n_{\ell_{1},...\ell_{k}}(M)}}{t^{k}\ell_{1}\cdots\ell_{k}}=\overline{\mathcal{T}}_{\ell}(t,\mathbf{t}).$
@@
Note that the automorphism group of any BMS map is trivial. This follows from the fact that every connected component of a BMS map will have at least one distinguished point. By the argument in Proposition 1, one can see that the generating functions of BMS maps have integer coefficients, that is, $\overline{\mathcal{T}}_{\ell}(t,\mathbf{t})\in\mathbb{N}[\mathbf{t}][[t]]$.
@@ 3.2 A review of ordinary map enumeration
Let’s very briefly review the standard methods used to enumerate ordinary maps, whose techniques will be useful when enumerating stuffed maps. Tutte originally constructed a recursive bijection for ordinary maps by removing the rooted edge of a boundary and reassigning new root(s) [tutte1968enumeration]. The resulting recursive formula for generating functions is
@@
For a detailed review, we refer the reader to Chapter 2 of [chapuy2014introduction]. Note that when all weights are zero, the recursion is for planar rooted one-faced maps, the solution of which is the Catalan numbers, i.e. the number of unicellular maps. By defining a generating function of $T_{\ell}$’s, one can then write the above equations as a quadratic equation. We define the formal generating function of maps with one boundary as
@@
Now assume that the number of 2-cells used to construct maps is finite, i.e. $|S|<\infty$, and that all 2-cells have boundaries of even length. Both $V$ and $R$ are polynomials in $x$, and by a theorem of Brown [brown1965existence] the discriminant is a polynomial that factorizes as
@@
The generating function $W(x)$ is much more simple to solve for in Zhukovsky space. For details, we refer the reader to Chapter 3.1 of [eynard2016counting].
@@
equivalent to what is known as the tree equation. It is called such because it is often derived by bijective methods involving decorated trees [cori1981planar, bouttier2004planar]. From this equation, many of the statistical properties and the critical behaviour of random maps can be studied [bouttier2003geodesic, marckert2007invariance]. Consider a map enumerated by $T^{\circ}_{2}$. If one zips closed the 2-gon boundary, the result is a map with no boundary but one rooted edge. Hence $T^{\circ}_{2}-t$ is the generating of rooted pointed maps, and via the above formula $2\gamma^{2}$ is also. The generating function $\gamma^{2}$ enumerates pointed maps with a distinguished edge. Via the BDFG bijection, this is the generating function of mobiles.
Further, all generating functions $T_{2\ell}$ can be computed in terms of $\gamma$ by expanding $T_{2\ell}=-\text{Res}_{x\rightarrow\infty}W(x)x^{2\ell}dx$ in Zhukovsky coordinates. The result found in [eynard2016counting] is

### v2

# Enumerating planar stuffed maps as hypertrees of mobiles

@@ Abstract
A planar stuffed map is an embedding of a graph into the 2-sphere $S^{2}$, considered up to orientation-preserving homeomorphisms, such that the complement of the graph is a collection of disjoint topologically connected components that are each homeomorphic to $S^{2}$ with multiple boundaries. This is a generalization of planar maps whose complement of the graph is a collection of disjoint topologically connected components that are each homeomorphic to a disc. The main goal of this work is to construct a bijection between bipartite planar stuffed maps and collections of integer-labelled trees connected by hyperedges such that they form a hypertree, called hypermobiles. This bijection directly generalizes the Bouttier-Di Franceso-Guitter bijection between bipartite planar maps and mobiles. Additionally, we show that the generating functions of these trees of mobiles satisfy both an algebraic equation, generalizing the case of ordinary planar maps, and a new functional equation. As an example, we explicitly enumerate a class of stuffed quadrangulations.
@@ 1 Introduction
Planar maps were first studied and enumerated by Tutte in the 1960’s [29] and have since appeared in statistical mechanics, quantum gravity, and probability theory. For a detailed history see [28]. One key insight discovered was that planar maps are in bijection with decorated trees [11, 27] from which enumerative formulae could be derived. Of particular relevance to this work is the bijection of Bouttier-Di Franceso-Guitter (BDFG) [7], which is between planar maps and plane trees with integer-labelled vertices called mobiles. Such bijections have since been generalized beyond planar maps [9, 25]. In addition to enumeration, these bijections gave insight into the discrete geometry of maps which has found applications in theoretical physics and probability theory [6, 21].
@@
As the reader will see, stuffed maps have many interesting combinatorial properties, and their enumeration is much more complicated than ordinary maps. Stuffed maps were first introduced in [5], where it was shown that stuffed maps of any genus with boundaries could be recursively enumerated using a process called blobbed topological recursion. This work was largely motivated by the study of multi-tracial matrix models. As models of quantum gravity, the multi-tracial terms in the potential of such matrix models contribute “wormholes" in the two-dimensional universes of their Feynman diagrams i.e. stuffed maps. In particular, an interior of a 2-cell with multiple boundaries corresponds to a wormhole and a path-connected component corresponds to universes in the physics literature [12, 2]. Such matrix models are also of interest in what is known as the duality of Liouville quantum gravity [19, 13, 26]. Other applications include the $O(n)$ model which involves ordinary maps decorated with self-avoiding loop configurations which can be shown to be equivalent to stuffed maps as well as maps with intersecting loops.
We hope that this paper will lay the foundations for studying the geometric and statistical properties of stuffed maps. Stuffed maps have yet to be fully studied from the perspective of random maps. Recently, much work has been done to connect the Gromov-Hausdorf limit of random metric spaces to random surfaces from two-dimensional quantum gravity. Researchers are looking for new random surfaces, and many types of maps have been shown to converge to such spaces [23]. It is believed, based on critical exponent analysis, that stuffed maps will not fall into the same class as ordinary maps, and hence might be related to new types of random surfaces. One can enumerate stuffed maps by computing formal multi-tracial matrix integrals, and in the physics literature such integrals are conjectured to have a continuum limit with "baby-universes" [20, 1]. Certain coloured stuffed maps appear also in the study of tensor models [4]. Establishing bijections with trees could be used as a first step in establishing a rigorous construction of a random surfaces with such features, as has been done for many types of ordinary maps [21, 22, 3, 15].
@@ 2.2 Mobiles
  • For any white vertex labelled $n$ that is attached to a black vertex, any adjacent clockwise white vertex attached to the same black vertex must be labelled at least n1n-1.
@@
The BDFG bijeciton is between pointed rooted bipartite planar maps and rooted mobiles [7], which generalized the bijection in [27] between quadrangulations and labelled trees.
@@
    For every corner of a white vertex in $m$, if the label of the vertex is $n_{\text{min}}(m)$ then connect it to the white vertex added for that mobile in the previous step. Otherwise, if the label is $n\not=n_{\text{min}}(m)$, then connect it to the next white vertex labelled n1n-1 in a clockwise direction around the contour. If this corner is the corner to the left of a root, the new edge is rooted pointing away for the corner.
@@ 2.3 Hypermobiles
  • A set of mobiles $\mathcal{M}$ with a unique mobile with the lowest labelled vertex, called the origin mobile.
@@ 2.4 Branch-pointed stuffed maps
As mentioned in the Introduction, one motivation to study stuffed maps is to establish convergence, in some sense, as a random discrete metric space to a random continuous metric space. All methods known to the author require a bijection to some path-connected tree-like object. However, one distinguishing feature of stuffed maps is that they may have distinct path-connected components connected by branches, which makes it not always possible to label vertices with their graph distances from the source vertex. To facilitate such bijections, we will slightly change the type of stuffed map studied in this work by decorating them and restricting their 2-cells. Define a branch-pointed stuffed (BPS) map as a stuffed map such that each boundary of a multi-boundary 2-cells has a distinguished vertex called a spurious point. There is an additional distinguished point that does not coincide with a spurious point. The source cannot coincide with a spurious point. One can define the following set of BPS maps analogously to our definition of $\mathbb{M}_{\ell}(S;v)$ for stuffed maps: denote it $\overline{\mathbb{M}}_{\ell}(S;v)$.
Working with rooted BPS maps allows one to give each vertex a positive integer label, which is key to the bijection. Note that on the level of enumeration, computing the generating function of BPS maps suffices to enumerate stuffed maps and vice versa. For more details, see the Remark in Section 3.1.
@@ 2.5 A bijection with hypermobiles
In this section we will construct and prove the main bijection between planar BPS maps and hypermobiles. First, consider the mapping $\Phi$ that takes a rooted pointed BPS map $M\in\overline{\mathbb{M}}^{\circ}(S;v)$ to a hypergraph $\Phi(M)$, constructed as follows:
@@
Figure 4: Stages of the construction of a hypermobile from the BPS map seen in Figure 1 (b). Sub-figure (a) is a BPS map whose vertices are labelled with the distance of the source and whose edges next to gate vertices are connected by hyperedges. In sub-figure (b), edges are drawn according to the same rules as the BDFG bijection on each component. Sub-figure (c) is the resulting hypermobile after removing all the original edges and branches, and shifting labels.
@@
    For every corner of a white vertex in $m$, then connect it to the next vertex in a clockwise direction with n1n-1 in the contour. If this corner is the corner to the left of a root, the new edge is rooted pointing away for the corner.
@@
    Remove all the original edges and black vertices of $m$. Replace each hyperedge with a 2-cell whose boundaries correspond to the hyperedges’ vertices. The minimum labelled vertex on the origin mobile is taken to be the origin. Spurious points are taken to be the vertex in a counterclockwise direction along each boundary face of a gate vertex.
@@
The mapping $\Phi$ is equal to $\Psi^{-1}$ and is a bijection between rooted and pointed BPS maps $\overline{\mathbb{M}}^{\circ}(S;v)$ and hypermobiles $\mathbb{T}(S;v)$.
@@
First, we will show that $\Phi$ maps into $\mathbb{T}(S;v)$. Let $M\in\overline{\mathbb{M}}^{\circ}(S;v)$. It is clear that $\Phi(M)$ is a labelled hypergraph. On each component one has applied the BDFG bijection, so they correspond to mobiles. The resulting graph is a collection of mobiles connected by the occasional hyperedges. Consider a host graph $G$ formed from $\Phi(M)$ by replacing each hyperedge with a linear sequence of n1n-1 red edges connecting all black vertices in that hyperedge. If we can show that $G$ is a tree, then this suffices to show that $\Phi(M)$ is a hypertree. Suppose for purposes of contradiction that $G$ contains a loop. No loop can be formed with just ordinary edges since this would contradict that one of the mobiles is a tree. Thus, such a loop must contain at least one red edge. However, any such sequence of edges that forms a cycle would have to come from a sequence of branches and components that topologically forms a handle in $M$, which contradicts the planarity of $M$. Thus, $G$ has no cycles.
What remains is to show that $G$ is connected. By construction, recall that there are $F(M)$ many black vertices, and V(M)K(M)V(M)-K(M) many white vertices. The number of edges between white and black vertices is $E(M)$, since for each existing edge precisely one edge is created. The number of red edges is $\sum_{\ell}|\ell|$, where the sum is over all distinct linear sequences $\ell$ of red edges in $G$. Note that any such sequence corresponds to the number of boundaries of the corresponding branch in $M$ minus one, i.e.
@@
where $B_{i}(M)$ denotes the number of branches in $M$ with $i$ boundaries. Recall from Section 2.1 the associated hypertree to $M$ and note that it must have a host tree. As a tree, this implies that by its construction B(M)=K(M)1B(M)=K(M)-1 and the degree sum formula can be written as $\sum_{i=2}iB_{i}(M)=2B(M)$. Putting this all together:
@@ 3.1 Generating functions
From generalizing the well-known Euler characteristic argument (see Theorem 1.2.1 of [14] for example) we will show that $\mathbb{M}_{\ell}(S;v)$ is a finite set, implying that $\mathcal{T}_{\ell}(t,\mathbf{t})$ is a well-defined formal series.
@@
Let $\overline{\mathcal{T}}_{\ell}(t,\mathbf{t})$ denote the generating function above but for BPS maps. The difference between the generating functions of stuffed maps and branch-rooted stuffed maps is a symmetry factor of $\ell_{1}\cdots\ell_{k}$ for every 2-cell with boundaries of lengths $\ell_{1},...,\ell_{k}$ and a factor of $t$ for each boundary, i.e.
[…]
$=\overline{\mathcal{T}}_{\ell}(t,\mathbf{t}).$
@@
Note that the automorphism group of any BPS map is trivial. This follows from the fact that every connected component of a BPS map will have at least one distinguished point. By the argument in Proposition 1, one can see that the generating functions of BPS maps have integer coefficients, that is, $\overline{\mathcal{T}}_{\ell}(t,\mathbf{t})\in\mathbb{N}[\mathbf{t}][[t]]$.
@@ 3.2 A review of ordinary map enumeration
Let’s very briefly review the standard methods used to enumerate ordinary maps, whose techniques will be useful when enumerating stuffed maps. Tutte originally constructed a recursive bijection for ordinary maps by removing the rooted edge of a boundary and reassigning new root(s) [29]. The resulting recursive formula for generating functions is
@@
For a detailed review, we refer the reader to Chapter 2 of [10]. Note that when all weights are zero, the recursion is for planar rooted one-faced maps, the solution of which is the Catalan numbers, i.e. the number of unicellular maps. By defining a generating function of $T_{\ell}$’s, one can then write the above equations as a quadratic equation. We define the formal generating function of maps with one boundary as
@@
Now assume that the number of 2-cells used to construct maps is finite, i.e. $|S|<\infty$, and that all 2-cells have boundaries of even length. Both $V$ and $R$ are polynomials in $x$, and by a theorem of Brown [8] the discriminant is a polynomial that factorizes as
@@
The generating function $W(x)$ is much more simple to solve for in Zhukovsky space. For details, we refer the reader to Chapter 3.1 of [14].
@@
equivalent to what is known as the tree equation. It is called such because it is often derived by bijective methods involving decorated trees [11, 7]. From this equation, many of the statistical properties and the critical behaviour of random maps can be studied [6, 24]. Consider a map enumerated by $T^{\circ}_{2}$. If one zips closed the 2-gon boundary, the result is a map with no boundary but one rooted edge. Hence $T^{\circ}_{2}-t$ is the generating of rooted pointed maps, and via the above formula $2\gamma^{2}$ is also. The generating function $\gamma^{2}$ enumerates pointed maps with a distinguished edge. Via the BDFG bijection, this is the generating function of mobiles.
Further, all generating functions $T_{2\ell}$ can be computed in terms of $\gamma$ by expanding $T_{2\ell}=-\text{Res}_{x\rightarrow\infty}W(x)x^{2\ell}dx$ in Zhukovsky coordinates. The result found in [14] is